OpenStudy (anonymous):

Fan and Medal! Help please! Carbon-14 decays at a constant rate, so it can be used to determine the age of fossils. In particular, if the original amount of Carbon-14 present is A0, then A(t) = A0e^(-kt) can be used find the amount of amount of Carbon-14 remaining after t years. Given that the half-life of Carbon-14 is 5,730 years, what is the value of the decay constant k to 5 decimal places?

OpenStudy (anonymous):

1/2mo=moxe^(-k5730) 0.5=e^(-k5730) ln0.5=-k5730 -0.693=-k5730 k=-0.693/-5730 k=0.000120